OpenStudy (anonymous):

A wire is first bent into the shape of a rectangle with width 4in and length 5in. Then the wire is unbent and reshaped into a triangle. If each side of the triangle has equal length, what is the length.

OpenStudy (anonymous):

find the area

OpenStudy (anonymous):

this is all new to me thats not really helping much

OpenStudy (anonymous):

 perimeter= 2(l+w) =2(4+5) =18 3x=18 x=18/3 x=6 so length=6